A 1.2 kg ball drops vertically onto afloor, hitting with a speed of 20 m/s.It rebounds with an initial speed of 10 m/s.
(a) What impulse acts on the ball during thecontact?
kg·m/s

(b) If the ball is in contact with the floor for 0.020 s, what is the average force exerted on thefloor?
N
Answer #1
(a) Impulse,
              I = F.Δt = Δp = m(v - u)
                  =1.2(10 - - 20) = 36kg.m/s
(b) Δt = 0.020 s
Average force,
             F = I/Δt = 36/0.02 = 1800 N